OpenStudy (anonymous):

Write the standard equation for the circle. center (10, –6), r = 6 A. (x + 10)2 + (y – 6)2 = 6 B. (x + 6)2 + (y – 10)2 = 36 C. (x – 10)2 + (y + 6)2 = 36 D. (x – 10)2 + (y + 6)2 = 6

OpenStudy (anonymous):

C. (x – 10)2 + (y + 6)2 = 36 standard form is (x-x1)^2 + (y-y1)^2=r^2

OpenStudy (anonymous):

thanks